OpenStudy (anonymous):

Please Help!! A bullet of mass 5 grams is fired with a velocity of 350 m/s from a 4 kg gun. Calculate: a) the recoil velocity of the gun b) the force on the bullet if the expanding gases act on the bullet for .006 s I already figured out the recoil velocity which is .0437 m/s thanks!

OpenStudy (anonymous):

oh ok so the expanding gases will keep acting on the bullet until it stops?

OpenStudy (anonymous):

a = .437/.006 = 72.833 m/s^2 ?

OpenStudy (anonymous):

F= .005 * -72.83 = -.3641 N why does it seem so little

OpenStudy (anonymous):

oh lol yeah me too..

OpenStudy (lena772):

So the acceleration is -7.28 m/s^2

OpenStudy (anonymous):

F= .005 * -7.28 = -.0364 ?

OpenStudy (anonymous):

should I be using the bullets velocity instead of the recoil? :S so confused

OpenStudy (anonymous):

yeah i did convert to kg

OpenStudy (lena772):

Ok I just did this whole problem wrong, I'm so sorry. I didn't realize that the bullet's velocity was given.

OpenStudy (lena772):

Okay so force equals mass x acceleration. We have the mass of the bullet, 5g. However, we do not have the acceleration. Can we find the acceleration? Yes, we can. The formula for acceleration is vf-vi/t vf=final velocity vi= initial velocity. Velocity is displacement over time. You have the final velocity of the bullet given, which is 350 m/s, before the bullet was fired, it had not been displaced, so the velocity was 0 m/s. The time is a given. It is 0.006 seconds. Now we will plug the numbers into the equation.

OpenStudy (anonymous):

a=58333 m/s^2

OpenStudy (anonymous):

350m/s / .006s = 58333 m/s^2

OpenStudy (anonymous):

it can't be smaller than 350 - the denominator is less than zero

OpenStudy (anonymous):

~less than 1

OpenStudy (lena772):

My calculator is giving me the wrong answer

OpenStudy (anonymous):

blasted technology!! ^_^

OpenStudy (lena772):

A = VF-VI/T A= 350m/s-0m/s/0.006s A=350/0.006 A=58333m/s2 Force = mxa Force = 0.005(convert the 5g to kg)*58333 Force = 291.67 N

OpenStudy (anonymous):

okay thanks guys :)))